A projectile is launched with an initial speed of 46.0 m/s at an angle of 35.0° above the horizontal. The projectile lands on a hillside 3.95 s later. Neglect air friction. (Assume that the +x-axis is to the right and the +y-axis is up along the page.)

(a) What is the projectile's velocity at the highest point of its trajectory?


1
Expert's answer
2021-08-19T14:23:13-0400

"v_0 = 46\frac{m}{s}"

"\alpha = 35\degree"

"t_1 = 3.95s"

"g= 9.8\frac{m}{s^2}"

"\text {Projectile rise time:}"

"t_p = \frac{v_0\sin\alpha}{g}= \frac{46*\sin35\degree}{9.8}=2,69s"

"t_p<t_1\text{ that is, the projectile reaches the top point of the trajectory}"

"\text{{ without colliding with the mountainside}}"

"\text{for }t_p:"

"v_y = 0;v_x= v_0\cos\alpha;"

"v = 46*\cos35\degree=37.68\frac{m}{s}"


"\text {Answer: }v =37.68\frac{m}{s}"